How many moles of H2O will be produced from 2.9 g of HCl reacting with Ca(OH)2?
1
Expert's answer
2020-05-01T14:22:23-0400

Ca(OH)2 + 2HCl = CaCl2 + 2H2O


Moles of H2O = moles of HCl = 2.9/36.5 = 0.0798 mol